x^3-3x^2+3x-9=0
I got x=3; x=3i; x=-3i...how do I plug answers back in to check solution?

Answer by venugopalramana(3286)

LET Y = x^3-3x^2+3x-9=0
I got x=3; x=3i; x=-3i...how do I plug answers back in to check solution?
HOW DID YOU GET THOSE ANSWERS?...ANY WAY ,YOU HAVE TO JUST SUBSTITUTE THOSE VALUES FOR X AND CHECK
1...X=3...Y=3^3-3*3^2+3*3-9=27-27+9-9=0...OK..
2...X=3i..Y=(3i)^3-3*(3i)^2+3*(3i)-9=27*(i^2)(i)-27(i)^2+9i-9=-27i+27+9i-9=-18i+18...
THIS IS NOT ZERO HENCE THIS IS NOT A ROOT
3...X=-3i..Y=(-3i)^3-3*(-3i)^2+3(-3i)-9=-27(i^2)(i)-27(i^2)-9i-9=27i+27-9i-9=18i+18...
THIS IS NOT ZERO ..HENCE THIS IS NOT A ROOT..
THE CORRECT ANSWERS ARE
X=3
X=i*SQRT(3)
X=-i*SQRT(3)
YOU CAN CHECK BY SUBSTITUTION AS SHOWN ABOVE